OpenStudy (anonymous):

The dimensons of a rectangle are such that its length is 5 inches more than its width. If the lenght were doubled and if the width were decreased by 2 inches, the area would be increased by 112 inches squared. What are the length and width of the rectangle?

OpenStudy (anonymous):

Hi Welcome to Openstudy! You are given: width= x length=5 + x (length is 5 inches more than its width) Then it says: If the lenght were doubled : 2(5 + x) the width were decreased by 2 inches: (x-2) therefore you'll have: length: 2(5+x-2)=2(x+3) = 2x + 6 width= x-2 Then area would be increased by 112 inches squared. Area of a rectangle is \(A=l \times w\) so next we have to do is combine these formulas or simplify them.. do you understand so far? Original Area without conditions is: \(A= (x+5)(x)\) New Area with the condition is: \(A+112= (x+3)(x-2)\)

OpenStudy (anonymous):

*I mean combine and simplify \(A=(x+5)(x)\ and\ A+112=(x+3)(x-2)\)

OpenStudy (anonymous):

@Missy1101 still there?

OpenStudy (anonymous):

so let's simplify the equations: \(A=(x+5)(x)=x^2+5x\) \(A+112=(x+3)(x-2)=x^2+x-6\\A=x^2+x-6-112\\A=x^2+x-118\) combine these two equations: \(x^2+5x = x^2+x-118\) since A is common between these equations.. Now it's your turn to continue.. can you simplify: \(x^2+5x = x^2+x-118\) ? and calculate for the value of x (width) I'll wait for your response :)